OpenStudy (fionacndg):

How can 1/5x – 2 = y and 1/3x +8 = y be set up as a system of equations?

OpenStudy (anonymous):

1/5x-2 = 1/3x + 8 You can set each equal to y and you will have two equations: y = 1/5x - 2 y = 1/3x + 8 multiply the first equation throughout by 5 to get rid of the fraction. multiply the first equation throughout by 3 to get rid of the fraction. In each equation, keep the constant on the right hand side and bring everything else to the left hand side.

OpenStudy (fionacndg):

Okay... let me process this!

OpenStudy (anonymous):

k

OpenStudy (fionacndg):

so would y = 1/5x - 2 --> y= 5x-10

OpenStudy (fionacndg):

@pielover123 so the first equation = -10 right? so what am i doing wrong or what do i do from here

OpenStudy (anonymous):

your doing it right

OpenStudy (fionacndg):

okay and the second equals x+24

OpenStudy (anonymous):

Yes sir!

OpenStudy (fionacndg):

I have -34 as my answer. what do i do now?

OpenStudy (fionacndg):

here is an attachment for possible answers I have.

OpenStudy (fionacndg):

@pielover123

OpenStudy (anonymous):

it would be the second one if i'm correct

OpenStudy (fionacndg):

Okay. is there a simple way I can check that?

OpenStudy (anonymous):

Not that i know of

OpenStudy (fionacndg):

Okay. thank you!!!!! I am hoping that it is correct!

OpenStudy (fionacndg):

unfortunately that wasn't the answer.

OpenStudy (anonymous):

Im sorry

OpenStudy (fionacndg):

Not like I could do any better! Thank you for your time! have awesome day!

OpenStudy (anonymous):

You to cya